# A MIND for Reasoning: Meta-learning for In-context Deduction

### v1

@@ 9 Limitations
Despite demonstrating meaningful progress in enhancing the deductive reasoning capabilities of language models through the MIND approach, this study has several limitations that future research could address.
@@
The evaluation primarily targets small to mid-sized language models (1.5B to 7B parameters), largely due to computational constraints. This focus leaves open the question of whether the observed improvements from MIND generalize to larger-scale models.
@@
The gains in reasoning ability achieved by MIND come with associated costs. The meta-learning strategy adopted involves incorporating multiple study examples into the input context during fine-tuning. This leads to longer input sequences, which in turn increase memory usage and computational demands compared to standard fine-tuning approaches.
@@
This work is constrained to the syllogistic fragment of first-order logic. Future research should investigate whether our conclusions extend to more expressive logical systems or to real-world scenarios where reasoning tasks are less structured. However, syllogistic logic is a restricted domain that allows for precise control over variables such as the type of inference considered, inference length, and the structure of knowledge bases. In the context of this study, it serves as a valuable testbed for investigating logical generalization in LLMs.
@@ Appendix A Formal Semantics and Syllogistic Inference Patterns
In this section, we formally define the semantics of syllogistic logic by translating syllogistic formulas into first-order logic. We also specify a consistent set of such formulas and formalize a valid inference within this framework. Let $\mathcal{A}=\{a,b,c,\ldots\}$ be a set of atomic terms, and let =$\mathcal{R}=\{R,S,T,\ldots\}$ be a set of unary relational symbols. We bijectively assign to every atomic term $a\in\mathcal{A}$ a relational symbol $R_{a}\in\mathcal{R}$, and interpret syllogistic formulas as first-order logic sentences: $Aab$ as $\forall x\,[R_{a}(x)\to R_{b}(x)]$, $Eab$ as $\forall x\,[R_{a}(x)\to\neg R_{b}(x)]$, $Iab$ as $\exists x\,[R_{a}(x)\land R_{b}(x)]$, and $Oab$ as $\exists x\,[R_{a}(x)\land\neg R_{b}(x)]$. We say that a set $\mathcal{F}$ of syllogistic formulas is consistent if there exists a structure $M$ in signature $\mathcal{R}$ such that every relation $R^{M}$ is non-empty, and the interpretation of every sentence in $\mathcal{F}$ holds in $M$, denoted by $M\vDash\mathcal{F}$. For a syllogistic formula $F$, the pair $(\mathcal{F},F)$ is an inference, denoted by $\mathcal{F}\vDash F$, if $M\vDash\{F\}$, whenever $M\vDash\mathcal{F}$ for a structure $M$ in signature $\mathcal{R}$.
@@ 7 $\{Aa-b,Ac-d,Iac\}\vDash Ibd$
Table 4: Syllogistic inference types. Each row shows a distinct logical inference pattern. Notation follows traditional categorical logic: $Aab$ denotes a universal affirmative ("All $a$ are $b$"), $Eab$ a universal negative ("No $a$ are $b$"), $Iac$ a existential affirmative ("Some $a$ are $c$"), and $Oad$ a existential negative ("Some $a$ are not $d$"). Formulas of the form $Aa-b$ denote a sequence of $n$ $A$-formulas relating $a$ and $b$.
B.1 $\mathcal{KB}$s’ Generation
Knowledge bases can be modeled as edge-labeled graphs, in which nodes correspond to atomic terms and edges are labeled with quantifiers. Our graph generation algorithm comprises two principal stages: (1) We first construct all A-chains of the knowledge base, which is used as its structural backbone, by generating disjoint trees—directed acyclic graphs that ensure a unique path exists between any pair of nodes. (2) Subsequently, we incorporate additional label edges corresponding to $E$, $I$, and $O$ formulas, while maintaining the overall consistency of the knowledge base.
@@ B.2 $\mathcal{KB}$s’ Visualization
Figure 19 illustrates a type 2 inference, characterized by a conclusion in the form of a universal affirmative ($A$-formula). The premises consist of a single sequence of $A$-formulas. This represents the most elementary form of syllogistic inference, whose structural pattern is embedded within all other types. Inferences of types 1, 3, and 5, which yield particular negative conclusions ($O$-formulas), are presented in Figures 1820, and 22, respectively. Syllogisms corresponding to types 4 and 7, both concluding with particular affirmative statements ($I$-formulas), are shown in Figures 21 and 24. Finally, the type 6 inference, which concludes with a universal negative ($E$-formula), is depicted in Figure 23.
@@ C.1 Implementation Details
All experiments were conducted using the PyTorch and Hugging Face Transformers libraries. We used NVIDIA A100 80GB GPUs. Due to the relatively small size of the models used in the experiments, each fine-tuning run, both for MIND and the baseline, was able to fit on a single GPU. We estimate a total compute usage of approximately 500 GPU hours across all experiments. Additionally, GitHub Copilot was used as an assistant tool for parts of the project’s source code development.
@@ C.2 Fine-tuning Details
All models were fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al., 2022) with a rank $r=64$, alpha value $\alpha=128$, and dropout probability $p=0.05$. The adaptation was applied to all attention and linear weight matrices, excluding the embedding and unembedding layers. Baseline models were loaded in bfloat16 precision, while MIND fine-tuned models employed QLoRA (Dettmers et al., 2023) with 4-bit quantization to accommodate memory constraints from longer sequences. Despite the lower precision, the meta-learning models outperformed the baseline.
@@ C.3 Closed Source Models
Qwen-2.5 1.5B MIND 93.11 ± 0.61 93.15 ± 0.11 74.24 ± 1.07
Baseline 85.56 ± 1.24 83.34 ± 1.90 38.49 ± 1.06
Qwen-2.5 3B MIND 96.16 ± 0.44 96.09 ± 0.30 83.21 ± 1.19
Baseline 93.03 ± 1.15 91.49 ± 0.68 53.12 ± 2.03
Qwen-2.5 7B MIND 98.13 ± 0.98 98.03 ± 1.19 86.87 ± 0.31
Baseline 95.76 ± 1.10 94.89 ± 1.55 57.81 ± 2.17
Table 6: Lexical generalization. Accuracy (mean ± std) of MIND and Baseline models in core generalization as in the main paper (Core) and with novel unseen terms (Unseen Pseudowords, Unseen Constants).
@@
Qwen-2.5 1.5B MIND 76.67 ± 0.38 50.40 ± 3.45 45.81 ± 1.13
Baseline 55.14 ± 0.53 29.37 ± 1.85 30.22 ± 1.52
Qwen-2.5 3B MIND 84.68 ± 0.54 64.77 ± 0.73 53.95 ± 3.46
Baseline 66.51 ± 0.19 43.66 ± 1.93 43.67 ± 2.05
Qwen-2.5 7B MIND 88.01 ± 1.11 69.24 ± 9.79 60.90 ± 2.94
Baseline 68.54 ± 2.25 45.27 ± 0.95 43.94 ± 2.82
@@ D.1 Accuracies by Type and Length
In this section, we present the complete set of accuracies broken down by type and length for both MIND and baseline models, as well as closed source models.
MIND and baseline.
We report the average accuracy for each inference type and length combination in both the core and length generalization settings for the Qwen-2.5 models. Figures 7, 8, and 9 show the accuracies for core generalization for the 1.5B, 3B, and 7B models, respectively, in both the MIND and baseline settings. Figures 13, 14, and 15 show the accuracies for short to long generalization, while Figures 10, 11, and 12 show the accuracies for long to short generalization for the same models, again in both the MIND and baseline settings.
Across model sizes and approaches, the easiest types of inferences are type 2 and type 6. In contrast, types 1, 3, and 4 are typically the most challenging. A notable difference between the MIND and baseline models is that the latter consistently struggle with type 5 inferences, whereas the former show stronger performance. However, apart from type 5 inferences, MIND models generally perform better but still tend to struggle or excel in similar type and length combinations as the baseline models.
@@
Figures 16 and 17 show the accuracies for zero-shot and few-shot prompting of GPT-4o and o3-mini, respectively. Both models show substantial improvement in the few-shot setting. GPT-4o is the lowest-performing model according to Table 1, a result further supported by the detailed breakdown in this section. It consistently achieves high accuracy only on type 2 inferences, which are the easiest and rely primarily on simple transitivity. o3-mini struggles more with types 3 and 4. Additionally, a clear difference in performance on type 5 inferences is observed between the zero-shot and few-shot settings. This resembles the difference seen in Qwen-2.5 models between MIND and baseline. These results show that even pretrained models tend to struggle with the same types of syllogistic inferences as fine-tuned models, with a few exceptions, such as type 5 inferences.
@@
Qwen-2.5 7B MIND 17.86 2.80 80.36 3.32 75.00
Baseline 6.67 5.19 91.43 5.39 80.95
@@ D.2 Lexical Generalization
Table 6 presents the results of this lexical generalization experiment. Across all Qwen-2.5 model sizes (1.5B, 3B, 7B) and conditions, the MIND approach consistently yields higher accuracy than the baseline, with performance improving with model scale for both approaches. Notably, for both known and unseen pseudowords, performance is similar in both the MIND and baseline settings, that is, changing the pseudoword vocabulary has little impact on model performance.
In contrast, for the most challenging generalization setting—unseen constants—both approaches exhibit a significant drop in performance, but the performance gap between MIND and the baseline becomes more pronounced: MIND achieves 86.87% at 7B, compared to just 57.81% for the baseline.
@@ D.3 Generalization with Limited Data
Table 7 presents the performance of the models when trained in a low data regime, using only 100 examples for each combination of inference type and length. Consistent with the findings in Table 6 and Table 2, MIND significantly outperforms the baseline across all model sizes and evaluation metrics. For the core generalization performance, the MIND models achieve substantially higher accuracy (e.g., 88.01% for Qwen-2.5 7B MIND vs. 68.54% for baseline). Similarly, when evaluating generalization to shorter and longer inferences than seen during training, MIND models demonstrate a clear advantage.
@@
D.4 Additional Error Analysis
In this section, we present the additional error analysis results for Qwen-2.5 7B both in MIND and baseline setting on the core generalization experiment. Additionally, we also show the error analysis results for GPT-4o and o3-mini. The detailed breakdown of these errors is presented in Table 8.
MIND and baseline.
For the Qwen-2.5 7B model, MIND shows a higher percentage of non-minimal valid set of premises (NVM) errors (17.86%) compared to the baseline (6.67%) on core generalization. However, when these NVM errors occur, MIND includes fewer unnecessary premises on average (Avg. NVM of 2.80) than the baseline (Avg. NVM of 5.19). Conversely, the baseline model exhibits a higher proportion of errors due to missing necessary A premises (MAP) at 91.43%, with an average of 5.39 missing premises. This is higher than MIND, which has a MAP percentage of 80.36% and an average of 3.32 missing premises. Both methods show high rates of hallucinated premises (HP), with MIND at 75.00% and the baseline slightly higher at 80.95%. These results suggest that not only MIND has generally a higher core generalization performance than the baseline, but also that MIND errors tend to be closer to the correct set of premises.
@@ D.4 Additional Error Analysis
Figure 7: Accuracy of MIND (Top) and Baseline (Bottom) Qwen-2.5 1.5B on core generalization decomposed by inference type and length.
Figure 8: Accuracy of MIND (Top) and Baseline (Bottom) Qwen-2.5 3B on core generalization decomposed by inference type and length.
Figure 9: Accuracy of MIND (Top) and Baseline (Bottom) Qwen-2.5 7B on core generalization decomposed by inference type and length.
Figure 10: Accuracy of MIND (Left) and Baseline (Right) Qwen-2.5 1.5B on long to short generalization decomposed by inference type and length.
Figure 11: Accuracy of MIND (Left) and Baseline (Right) Qwen-2.5 3B on long to short generalization decomposed by inference type and length.
Figure 12: Accuracy of MIND (Left) and Baseline (Right) Qwen-2.5 7B on long to short generalization decomposed by inference type and length.
Figure 13: Accuracy of MIND (Left) and Baseline (Right) Qwen-2.5 1.5B on short to long generalization decomposed by inference type and length.
Figure 14: Accuracy of MIND (Left) and Baseline (Right) Qwen-2.5 3B on short to long generalization decomposed by inference type and length.
Figure 15: Accuracy of MIND (Left) and Baseline (Right) Qwen-2.5 7B on short to long generalization decomposed by inference type and length.
@@
Figure 18: Type 1 syllogistic inference on graphs. Visualization of a type 1 syllogistic inference using a graph representation of an example $\mathcal{KB}$, alongside the corresponding textual translation. In the graph (top), nodes represent predicates. Black edges indicate A-formulas (“All As are Bs”), blue edges indicate I-formulas (“Some As are Bs”), red edges indicate E-formulas (“No As are Bs”), and yellow edges indicate O-formulas (“Some As are not Bs”). The query hypothesis is represented by a dashed green edge, and the edges that prove the hypothesis are highlighted in green. The text translation illustrates how the abstract graph representation is converted into a text format suitable for LM processing by applying fixed templates that represent logical formulas.

### v2

@@ 9 Limitations
Despite demonstrating meaningful progress in enhancing the logical reasoning capabilities of language models through the ML approach, this study has several limitations that future research could address.
@@
The evaluation primarily targets small to mid-sized language models (1.5B to 7B parameters). This focus leaves open the question of whether the observed improvements from ML generalize to larger-scale models.
@@
The gains in reasoning ability achieved by ML come with associated costs. The meta-learning strategy adopted involves incorporating multiple study examples into the input context during fine-tuning. This leads to longer input sequences, which in turn increase memory usage and computational demands compared to standard fine-tuning approaches.
Episode construction variability.
In this work, we fixed several meta-learning parameters that could be further explored in future studies. First, we set the number of study examples in each episode to three, primarily due to the memory constraints imposed by longer sequences in the ML setting. Second, our generalization experiments focused on length generalization, so the support set was restricted to inferences of the same type as the query. Future work could examine alternative support set configurations, for example investigating inter-inference transfer, that is, whether in-context knowledge of one inference type can be leveraged to solve queries of a different type.
@@
This work is constrained to the syllogistic fragment of first-order logic. Future research should investigate whether our conclusions extend to more expressive logical systems or to real-world scenarios where reasoning tasks are less structured. The future work should also evaluate meta-learning trade-offs for such more complex logical systems. However, syllogistic logic is a restricted domain that allows for precise control over variables such as the type of inference considered, inference length, and the structure of knowledge bases. In the context of this study, it serves as a valuable testbed for investigating logical generalization in LLMs.
@@ Appendix A Formal Semantics and Syllogistic Inference Patterns
In this section, we formally define the semantics of syllogistic logic by translating syllogistic formulas into first-order logic. We also specify a consistent set of such formulas and formalize a valid inference within this framework. Let $\mathcal{A}=\{a,b,c,\ldots\}$ be a set of atomic terms, and let $\mathcal{R}=\{R,S,T,\ldots\}$ be a set of unary relational symbols. We bijectively assign to every atomic term $a\in\mathcal{A}$ a relational symbol $R_{a}\in\mathcal{R}$, and interpret syllogistic formulas as first-order logic sentences: $Aab$ as $\forall x\,[R_{a}(x)\to R_{b}(x)]$, $Eab$ as $\forall x\,[R_{a}(x)\to\neg R_{b}(x)]$, $Iab$ as $\exists x\,[R_{a}(x)\land R_{b}(x)]$, and $Oab$ as $\exists x\,[R_{a}(x)\land\neg R_{b}(x)]$. We say that a set $\mathcal{F}$ of syllogistic formulas is consistent if there exists a structure $M$ in signature $\mathcal{R}$ such that every relation $R^{M}$ is non-empty, and the interpretation of every sentence in $\mathcal{F}$ holds in $M$, denoted by $M\vDash\mathcal{F}$. For a syllogistic formula $F$, the pair $(\mathcal{F},F)$ is an inference, denoted by $\mathcal{F}\vDash F$, if $M\vDash\{F\}$, whenever $M\vDash\mathcal{F}$ for a structure $M$ in signature $\mathcal{R}$.
@@ 7 $\{Aa-b,Ac-d,Iac\}\vDash Ibd$
Table 4: Syllogistic inference types. Each row shows a distinct logical inference pattern. Notation follows traditional categorical logic: AabAab denotes a universal affirmative ("All $a$ are bb"), EabEab a universal negative ("No $a$ are bb"), IacIac a existential affirmative ("Some $a$ are cc"), and $Oad$ a existential negative ("Some $a$ are not dd"). Formulas of the form AabAa-b denote a sequence of $n$ AA-formulas relating $a$ and $b$.
@@
B.1 $\mathcal{KB}$s’ Generation
Knowledge bases can be modeled as edge-labeled graphs, in which nodes correspond to atomic terms and edges are labeled with quantifiers. Our graph generation algorithm comprises two principal stages: (1) We first construct all A-chains of the knowledge base, which is used as its structural backbone, by generating disjoint trees—directed acyclic graphs that ensure a unique path exists between any pair of nodes. (2) Subsequently, we incorporate additional label edges corresponding to $E$, II, and $O$ formulas, while maintaining the overall consistency of the knowledge base.
@@ B.2 $\mathcal{KB}$s’ Visualization
Figure 20 illustrates a type 2 inference, characterized by a conclusion in the form of a universal affirmative ($A$-formula). The premises consist of a single sequence of $A$-formulas. This represents the most elementary form of syllogistic inference, whose structural pattern is embedded within all other types. Inferences of types 1, 3, and 5, which yield particular negative conclusions ($O$-formulas), are presented in Figures 1921, and 23, respectively. Syllogisms corresponding to types 4 and 7, both concluding with particular affirmative statements ($I$-formulas), are shown in Figures 22 and 25. Finally, the type 6 inference, which concludes with a universal negative ($E$-formula), is depicted in Figure 24.
@@ C.1 Implementation Details
All experiments were conducted using the PyTorch and Hugging Face Transformers libraries. We used NVIDIA A100 80GB GPUs. Due to the relatively small size of the models used in the experiments, each fine-tuning run, both for ML and the baseline, was able to fit on a single GPU. We estimate a total compute usage of approximately 500 GPU hours across all experiments. Additionally, GitHub Copilot was used as an assistant tool for parts of the project’s source code development.
@@ C.2 Fine-tuning Details
All models were fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al., 2022) with a rank $r=64$, alpha value $\alpha=128$, and dropout probability $p=0.05$. The adaptation was applied to all attention and linear weight matrices, excluding the embedding and unembedding layers. Baseline models were loaded in bfloat16 precision, while ML fine-tuned models employed QLoRA (Dettmers et al., 2023) with 4-bit quantization to accommodate memory constraints from longer sequences. Despite the lower precision, the meta-learning models outperformed the baseline.
@@ C.3 Closed Source Models
Qwen-2.5 1.5B ML 93.11 ± 0.61 93.15 ± 0.11 74.24 ± 1.07
$\mathrm{Baseline}_{-S}$ 85.56 ± 1.24 83.34 ± 1.90 38.49 ± 1.06
Qwen-2.5 3B ML 96.16 ± 0.44 96.09 ± 0.30 83.21 ± 1.19
$\mathrm{Baseline}_{-S}$ 93.03 ± 1.15 91.49 ± 0.68 53.12 ± 2.03
Qwen-2.5 7B ML 98.13 ± 0.98 98.03 ± 1.19 86.87 ± 0.31
$\mathrm{Baseline}_{-S}$ 95.76 ± 1.10 94.89 ± 1.55 57.81 ± 2.17
Table 6: Lexical generalization. Accuracy (mean ± std) of ML and Baseline models in core generalization as in the main paper (Core) and with novel unseen terms (Unseen Pseudowords, Unseen Constants).
@@
Qwen-2.5 1.5B ML 76.67 ± 0.38 50.40 ± 3.45 45.81 ± 1.13
$\mathrm{Baseline}_{-S}$ 55.14 ± 0.53 29.37 ± 1.85 30.22 ± 1.52
Qwen-2.5 3B ML 84.68 ± 0.54 64.77 ± 0.73 53.95 ± 3.46
$\mathrm{Baseline}_{-S}$ 66.51 ± 0.19 43.66 ± 1.93 43.67 ± 2.05
Qwen-2.5 7B ML 88.01 ± 1.11 69.24 ± 9.79 60.90 ± 2.94
$\mathrm{Baseline}_{-S}$ 68.54 ± 2.25 45.27 ± 0.95 43.94 ± 2.82
@@ D.1 Accuracies by Type and Length
In this section, we present the complete set of accuracies broken down by type and length for both ML and baseline models, as well as closed source models.
ML and baseline.
We report the average accuracy for each inference type and length combination in both the core and length generalization settings for the Qwen-2.5 models. Figures 8, 9, and 10 show the accuracies for core generalization for the 1.5B, 3B, and 7B models, respectively, in both the ML and baseline settings. Figures 14, 15, and 16 show the accuracies for short to long generalization, while Figures 11, 12, and 13 show the accuracies for long to short generalization for the same models, again in both the ML and baseline settings.
Across model sizes and approaches, the easiest types of inferences are type 2 and type 6. In contrast, types 1, 3, and 4 are typically the most challenging. A notable difference between the ML and baseline models is that the latter consistently struggle with type 5 inferences, whereas the former show stronger performance. However, apart from type 5 inferences, ML models generally perform better but still tend to struggle or excel in similar type and length combinations as the baseline models.
@@
Figures 17 and 18 show the accuracies for zero-shot and few-shot prompting of GPT-4o and o3-mini, respectively. Both models show substantial improvement in the few-shot setting. GPT-4o is the lowest-performing model according to Table 1, a result further supported by the detailed breakdown in this section. It consistently achieves high accuracy only on type 2 inferences, which are the easiest and rely primarily on simple transitivity. o3-mini struggles more with types 3 and 4. Additionally, a clear difference in performance on type 5 inferences is observed between the zero-shot and few-shot settings. This resembles the difference seen in Qwen-2.5 models between ML and baseline. These results show that even pretrained models tend to struggle with the same types of syllogistic inferences as fine-tuned models, with a few exceptions, such as type 5 inferences.
@@
Qwen-2.5 7B ML 17.86 2.80 80.36 3.32 75.00
$\mathrm{Baseline}_{-S}$ 6.67 5.19 91.43 5.39 80.95
@@ D.2 Lexical Generalization
Table 6 presents the results of this lexical generalization experiment. Across all Qwen-2.5 model sizes (1.5B, 3B, 7B) and conditions, the ML approach consistently yields higher accuracy than the baseline, with performance improving with model scale for both approaches. Notably, for both known and unseen pseudowords, performance is similar in both the ML and baseline settings, that is, changing the pseudoword vocabulary has little impact on model performance.
In contrast, for the most challenging generalization setting—unseen constants—both approaches exhibit a significant drop in performance, but the performance gap between ML and the baseline becomes more pronounced: ML achieves 86.87% at 7B, compared to just 57.81% for the baseline.
@@ D.3 Generalization with Limited Data
Table 7 presents the performance of the models when trained in a low data regime, using only 100 examples for each combination of inference type and length. Consistent with the findings in Table 6 and Table 2, ML significantly outperforms the baseline across all model sizes and evaluation metrics. For the core generalization performance, the ML models achieve substantially higher accuracy (e.g., 88.01% for Qwen-2.5 7B ML vs. 68.54% for baseline). Similarly, when evaluating generalization to shorter and longer inferences than seen during training, ML models demonstrate a clear advantage.
@@
D.4 Providing Study Examples to the Baseline at Inference Time
Table 6 presents a direct comparison between the two baseline variants: $\mathrm{Baseline}_{-S}$ (without study examples at inference time) and $\mathrm{Baseline}_{+S}$ (with study examples). In the main paper, we reported only $\mathrm{Baseline}_{-S}$ results, as it consistently performed better—or at least on par on average—across the settings highlighted there. Here, we provide a more detailed analysis.
Overall, both baseline variants underperform relative to the meta-learning approach, confirming that access to structured meta-learning episodes is crucial for systematic generalization. However, incorporating study examples into the baseline input does confer benefits in specific scenarios. In particular, $\mathrm{Baseline}_{+S}$ can outperform $\mathrm{Baseline}_{-S}$ in the low-data regime. This is likely because baseline models trained with limited data benefit from both training on task examples and leveraging their previously developed in-context learning capabilities, enabling better generalization from demonstrations of the inference task. For example, Qwen-2.5 3B achieves 75.80% (with study examples) versus 66.51% (without). A similar pattern emerges in the lexical generalization setting: $\mathrm{Baseline}_{+S}$ exhibits stronger robustness to unseen constants, which represent the largest distributional shift in vocabulary, again suggesting that study examples help the model generalize in-context from the learned vocabulary to novel symbols. For instance, Qwen-2.5 7B improves from 57.81% (S-S) to 74.28% (+S+S) when evaluated on unseen constants.
In the length generalization case, the effect of study examples depends critically on their alignment with test inference lengths. Specifically, disaligned study examples consistently reduce performance relative to $\mathrm{Baseline}_{-S}$. For instance, in the Short $\rightarrow$ Long task for Qwen-2.5 1.5B, disaligned examples yield 51.49% versus 63.53% without study examples, indicating that irrelevant in-context examples can confuse the model, particularly for longer inference lengths. Conversely, aligned study examples improve performance over disaligned ones, demonstrating that the baseline can exploit in-context cues when examples match the target length distribution. However, even with aligned examples, $\mathrm{Baseline}_{+S}$ rarely surpasses $\mathrm{Baseline}_{-S}$ for larger models (e.g., Qwen-2.5 7B, Short $\rightarrow$ Long: 79.90% aligned versus 80.76% without examples).
Since $\mathrm{Baseline}_{-S}$ tends to perform better in core generalization tasks, and the benefits of study examples at test time are inconsistent in length generalization scenarios, we reported $\mathrm{Baseline}_{-S}$ results in the main paper.
D.5 Additional Error Analysis
In this section, we present the additional error analysis results for Qwen-2.5 7B both in ML and baseline setting on the core generalization experiment. Additionally, we also show the error analysis results for GPT-4o and o3-mini. The detailed breakdown of these errors is presented in Table 8.
ML and baseline.
For the Qwen-2.5 7B model, ML shows a higher percentage of non-minimal valid set of premises (NVM) errors (17.86%) compared to the baseline (6.67%) on core generalization. However, when these NVM errors occur, ML includes fewer unnecessary premises on average (Avg. NVM of 2.80) than the baseline (Avg. NVM of 5.19). Conversely, the baseline model exhibits a higher proportion of errors due to missing necessary A premises (MAP) at 91.43%, with an average of 5.39 missing premises. This is higher than ML, which has a MAP percentage of 80.36% and an average of 3.32 missing premises. Both methods show high rates of hallucinated premises (HP), with ML at 75.00% and the baseline slightly higher at 80.95%. These results suggest that not only ML has generally a higher core generalization performance than the baseline, but also that ML errors tend to be closer to the correct set of premises.
@@ D.5 Additional Error Analysis
Model Method All Short Long
Core
Qwen-2.5 1.5B $\mathrm{Baseline}_{-S}$ 85.56 ± 1.24 91.42 ± 0.82 80.56 ± 1.78
$\mathrm{Baseline}_{+S}$ 77.89 ± 3.13 85.93 ± 2.17 70.58 ± 4.45
Qwen-2.5 3B $\mathrm{Baseline}_{-S}$ 93.03 ± 1.15 95.34 ± 1.18 90.92 ± 1.27
$\mathrm{Baseline}_{+S}$ 87.18 ± 4.32 91.98 ± 2.24 82.99 ± 6.41
Qwen-2.5 7B $\mathrm{Baseline}_{-S}$ 95.76 ± 1.10 97.27 ± 1.22 94.13 ± 0.90
$\mathrm{Baseline}_{+S}$ 92.91 ± 1.85 95.59 ± 1.06 90.29 ± 2.68
(a) Core generalization.
Model Method Short $\rightarrow$ Long Long $\rightarrow$ Short
Disaligned Aligned Disaligned Aligned
Qwen-2.5 1.5B $\mathrm{Baseline}_{-S}$ 63.53 ± 1.16 63.53 ± 1.16 56.67 ± 1.22 56.67 ± 1.22
$\mathrm{Baseline}_{+S}$ 51.49 ± 2.30 62.88 ± 4.59 60.38 ± 1.38 61.90 ± 2.08
Qwen-2.5 3B $\mathrm{Baseline}_{-S}$ 76.78 ± 1.63 76.78 ± 1.63 71.88 ± 1.49 71.88 ± 1.49
$\mathrm{Baseline}_{+S}$ 65.98 ± 4.02 69.55 ± 0.69 74.29 ± 1.95 74.84 ± 2.70
Qwen-2.5 7B $\mathrm{Baseline}_{-S}$ 80.76 ± 2.65 80.76 ± 2.65 71.08 ± 1.55 71.08 ± 1.55
$\mathrm{Baseline}_{+S}$ 75.34 ± 2.35 79.90 ± 1.53 73.88 ± 2.70 76.04 ± 1.95
(b) Generalization to unseen lengths.
[…]
Qwen-2.5 1.5B $\mathrm{Baseline}_{-S}$ 55.14 ± 0.53 29.37 ± 1.85 30.22 ± 1.52
$\mathrm{Baseline}_{+S}$ 66.18 ± 0.51 42.42 ± 1.48 28.35 ± 4.65
Qwen-2.5 3B $\mathrm{Baseline}_{-S}$ 66.51 ± 0.19 43.66 ± 1.93 43.67 ± 2.05
$\mathrm{Baseline}_{+S}$ 75.80 ± 0.15 58.50 ± 1.79 43.62 ± 4.13
Qwen-2.5 7B $\mathrm{Baseline}_{-S}$ 68.54 ± 2.25 45.27 ± 0.95 43.94 ± 2.82
$\mathrm{Baseline}_{+S}$ 74.69 ± 3.16 57.42 ± 3.07 45.60 ± 4.56
(c) Low-data regime.
Model Method Core Unseen Pseudowords Unseen Constants
Qwen-2.5 1.5B $\mathrm{Baseline}_{-S}$ 85.56 ± 1.24 83.34 ± 1.90 38.49 ± 1.06
$\mathrm{Baseline}_{+S}$ 77.89 ± 3.13 76.85 ± 2.85 54.22 ± 5.75
Qwen-2.5 3B $\mathrm{Baseline}_{-S}$ 93.03 ± 1.15 91.49 ± 0.68 53.12 ± 2.03
$\mathrm{Baseline}_{+S}$ 87.18 ± 4.32 85.94 ± 4.11 68.71 ± 6.74
Qwen-2.5 7B $\mathrm{Baseline}_{-S}$ 95.76 ± 1.10 94.89 ± 1.55 57.81 ± 2.17
$\mathrm{Baseline}_{+S}$ 92.91 ± 1.85 92.13 ± 1.93 74.28 ± 2.20
(d) Lexical generalization.
Figure 6: Baseline models with and without study examples. Each subfigure compares $\mathrm{Baseline}{-S}$ (without study examples at inference) and $\mathrm{Baseline}{+S}$ (with study examples). Results are shown for (a) core generalization, (b) generalization to unseen lengths, (c) low-data regime, and (d) lexical generalization. While $\mathrm{Baseline}{-S}$ generally performs better in fully trained settings, $\mathrm{Baseline}{+S}$ shows advantages in low-data and lexical generalization (unseen constants) scenarios, though both remain below meta-learning performance.
@@
Figure 8: Accuracy of ML (Top) and Baseline (Bottom) Qwen-2.5 1.5B on core generalization decomposed by inference type and length.
Figure 9: Accuracy of ML (Top) and Baseline (Bottom) Qwen-2.5 3B on core generalization decomposed by inference type and length.
Figure 10: Accuracy of ML (Top) and Baseline (Bottom) Qwen-2.5 7B on core generalization decomposed by inference type and length.
Figure 11: Accuracy of ML (Left) and Baseline (Right) Qwen-2.5 1.5B on long to short generalization decomposed by inference type and length.
Figure 12: Accuracy of ML (Left) and Baseline (Right) Qwen-2.5 3B on long to short generalization decomposed by inference type and length.
Figure 13: Accuracy of ML (Left) and Baseline (Right) Qwen-2.5 7B on long to short generalization decomposed by inference type and length.
Figure 14: Accuracy of ML (Left) and Baseline (Right) Qwen-2.5 1.5B on short to long generalization decomposed by inference type and length.
Figure 15: Accuracy of ML (Left) and Baseline (Right) Qwen-2.5 3B on short to long generalization decomposed by inference type and length.
Figure 16: Accuracy of ML (Left) and Baseline (Right) Qwen-2.5 7B on short to long generalization decomposed by inference type and length.
@@
Figure 19: Type 1 syllogistic inference on graphs. Visualization of a type 1 syllogistic inference using a graph representation of an example $\mathcal{KB}$, alongside the corresponding textual translation. In the graph (top), nodes represent predicates. Black edges indicate A-formulas (“All As are Bs”), blue edges indicate I-formulas (“Some As are Bs”), red edges indicate E-formulas (“No As are Bs”), and yellow edges indicate O-formulas (“Some As are not Bs”). The query hypothesis is represented by a dashed green edge, and the edges that prove the hypothesis are highlighted in green. The text translation illustrates how the abstract graph representation is converted into a text format suitable for LLM processing by applying fixed templates that represent logical formulas.